OpenStudy (anonymous):

What does tan^2(theta) mean?

OpenStudy (anonymous):

If\[\tan \theta = 1/3\] what is theta

hartnn (hartnn):

what exactly u wanna know about tan^2 theta ?? \(\tan \theta =1/3 \implies \theta =arctan(1/3)\) use calculator.

OpenStudy (anonymous):

tan theta means opposite/adjacent. So what does tan^2 theta mean? Square of opp/adj ?

zepdrix (zepdrix):

Yes.

hartnn (hartnn):

yeas, tan^2 theta is just square of tan theta \(\large \tan^2 \theta = (\frac{opposite}{adjacent})^2\)

OpenStudy (abb0t):

Remember: SOH: \[\sin = \frac{ opposite }{ hypotenuse }\] CAH \[\cos = \frac{ adjacent }{ hypotenuse }\] TOA \[\tan = \frac{ opposite }{ adjacent }\]

OpenStudy (kainui):

\[\tan^2(\theta)=(\tan \theta)^2\]

OpenStudy (kainui):

The reason we do this is because generally you don't write parenthesis, so it allows you to quickly and easily distinguish between the slightly ambiguous notation. \[\tan \theta^2\] and\[\tan^2 \theta\]

OpenStudy (anonymous):

if tan^2 = (opp/adj)^2 what does 2 tan mean?